OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (ivettef365):

you already have the y-int (0,-4) all you need to write the equation is the slope use the slope formula

OpenStudy (science0229):

Remember that slope for 2 points (a,b) and (c,d) is \[\frac{ c-a }{ d-b }\]

OpenStudy (science0229):

So the slope for (0,-4), (-2,2) is \[\frac{ 0--2 }{ -4-2 }=\frac{ 2 }{ -6 }=-\frac{ 1 }{ 3 }\]

OpenStudy (science0229):

Now, point (0,-4) is the y-intercept of the equation.

OpenStudy (science0229):

Just like ivettef365 said, you use the slope formula, which is \[y=mx+b\] where m is the slope and b is the y-intercept.

OpenStudy (science0229):

plug in -1/3 for m and 3 for b, and you'll get \[y=-\frac{ 1 }{ 3 }x-4\]

OpenStudy (ivettef365):

no, the slope is y2 - y1 ------ x2 - x1 2 - (-4) 6 ------- = ------ = -3 -2 -0 -2

OpenStudy (science0229):

Oops. You're right. My mistake. So the equation would be y=-3x-4

OpenStudy (ivettef365):

the equation is y = -3x -4

OpenStudy (science0229):

Thank you for correcting me.

OpenStudy (ivettef365):

your welcome

OpenStudy (anonymous):

Thanks guys!!(:

OpenStudy (ivettef365):

yw :)